write the standard form of an equation of the line that passes through each pair.

(-2,7-) (1,2)

Answer by stanbon(75887)

write the standard form of an equation of the line that passes through each pair.
(-2,7-) (1,2)
---------------------
slope = (2-7)/(1--2) = -5/3
----
Use the form y = mx + b to solve for "b":
2 = (-5/3)*1 + b
b = 11/3
---------
y = (-5/3)x + (11/3)
========================
Cheers,
Stan H.